OpenStudy (anonymous):

What type of sequence is this: 4, 6, 8, 10…?

OpenStudy (anonymous):

do you have to find the rule or something?

OpenStudy (anonymous):

no this is my only information! sorry!:(

OpenStudy (accessdenied):

This is the same question as the one before?

OpenStudy (anonymous):

Arithmetic sequence

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

the same question

OpenStudy (accessdenied):

http://openstudy.com/users/lovelychick13#/updates/53597a49e4b0fb65c3cab365 "What type of sequence is this: 4, 6, 8, 10…?" I think you posted the wrong one then, unless I've gone crazy. :p

OpenStudy (anonymous):

OH.... I DID!!! awww...

OpenStudy (anonymous):

What is the 20th term of the geometric sequence that starts with .02 and has a ratio of 2?

OpenStudy (accessdenied):

Our first term is .02. We multiply by our common ratio of 2 to get the second term. 0.02 * 2^1 = 2nd term For the 3rd term, multiply by 2 again: 0.02 * 2^2 = 3rd time Notice the pattern is thus, the n-th term is our initial term multiplied by the common ratio to the n-1 power: n-th term = 0.02 * 2^(n-1) We want the 20th, so just use that formula: 0.02 * 2^(20-1)

OpenStudy (anonymous):

ok...got .76

OpenStudy (accessdenied):

Just 0.76?

OpenStudy (anonymous):

ya why? is that wrong?

OpenStudy (accessdenied):

Oh I see, you multiplied: 0.02 * 2 * 19 The caret symbol I used (^) means an exponent: 0.02 * 2^(n-1) 0.02*2^(20-1) \( 0.02 * 2^{n-1} \) \(0.02 * 2^{20-1} \)

OpenStudy (anonymous):

wha does * mean?

OpenStudy (accessdenied):

Because for a geometric sequence, it is repeated multiplication. It is more convenient to write it as an exponent, though.

OpenStudy (accessdenied):

multiplication, like x or a dot. Sorry, I use a notation that I find more convenient to me. The same thing with a dot or multiplication sign is: \( 0.02 \times 2^{n-1} \) \( 0.02 \cdot 2^{n-1} \)

OpenStudy (anonymous):

I see! ok so my calculator cant do that...

OpenStudy (accessdenied):

What kind of calculator is it? Usually it is a key like \(x^y\) or if it is graphing calculator, it might just be a caret symbol ^

OpenStudy (anonymous):

oh I have it!! ok .04

OpenStudy (accessdenied):

I think that's only 0.02 * 2 If you use the key, you type your exponent afterwards so that your operations would be: 0.02 (multiply) 2 (exponent) (19)

OpenStudy (anonymous):

;( ummm....

OpenStudy (accessdenied):

Google calculator gives the answer 10485.76 https://www.google.com/#q=0.02*2%5E19 (The 0.76 confused me earlier from that)

OpenStudy (anonymous):

thanx!!!!!!!

OpenStudy (accessdenied):

Glad to help! :)